OpenStudy (anonymous):

I need help with this question!! It asks: Which of the following shows the correct steps to find the value of 8 to the power of 1 over 3?

OpenStudy (mathstudent55):

Can you show the rest of the problem?

OpenStudy (anonymous):

\[(\sqrt[3]{8?})^{1}\] Can you find the cubed root of 8?

OpenStudy (mathstudent55):

\(\huge 8^\frac{1}{3} = (2^3) ^\frac{1}{3} = 2^{3 \cdot \frac{1}{3} } = 2^1 = 2\)